OpenStudy (anonymous):

Simplify (4r^-2)*(2r^-2) Please show your workings.

OpenStudy (anonymous):

8r^-4 4*2 -2+(-2)

OpenStudy (anonymous):

My book gives the answer 8/r

OpenStudy (anonymous):

Chrissy, I'm pretty sure your book's answer is wrong, and tutordiffeq is right.

OpenStudy (anonymous):

4r^-2 = 4 / r^2 2r^-2 = 2 / r^2 so we have (4 / r^2) * (2 / r^2) = 8 / r^4

OpenStudy (anonymous):

Check to make sure you typed the question right.